Solve the following equation
sin(y')=0
1
Expert's answer
2020-08-27T13:51:43-0400

sin (y'(x)) = 0"\implies" The values for which sine function gives 0 value are n"\pi" , where n is an integer. So, here we have y'(x)= n"\pi"
=> y(x) = "\int" n"\pi" dx + C , where C is an integration constant.=> y(x) = n"\pi" x + C , which is the required answer to the given differential equation.